OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#11253 — RBX <> TH2
Scheduled Maintenance Report for Network & Infrastructure
Completed
Maintenance is taking place on the amplifier site in Nogent on the fibre cable between Roubaix and Paris - Telehouse 2.

The power equaliser is not operating correctly. We are going to reboot it and possibly replace it.

No impact on traffic is expected. The protection in GSW will take over.

Update(s):

Date: 2014-07-24 14:38:38 UTC
Intervention is complete. The TAC has validated the implementation of the new card.

Circuits are back up on the primary pathway.

Date: 2014-07-24 13:08:01 UTC
Th card has been replaced and the OCG cable between Roubaix and TH2 is back up. However, traffic will continue to be routed via GSW.

We will control with the TAC if everything goes OK.

Date: 2014-07-24 13:03:30 UTC
Circuit protection is currently going via GSW.

Date: 2014-07-24 13:02:30 UTC
We're replacing the equaliser.
Posted Jul 24, 2014 - 10:46 UTC